Physical Chemistry 14C.1(a)
14C. 1(a): The duration of a 90 degree or 180 degree pulse depends on the strength of the B1 field. If a 180 degree pulse requires 12.5 ms, what is the strength of the B1 field ? How long would the corresponding 90 degree pulse require?
Exercises 14C1(a) The duration of a 90° or 180° pulse depends on the strength of the feld. If a 180° pulse requires 12.5 us, what is the strength of the B, field? How long would the corresponding 90° pulse require?
Exercises 14C.1 1la) The duration of a 90° or 180° pulse depends on the strength of the B field if a 180° pulse requires 12.5 us, what is the strength of the field? How long would the corresponding 90 pulse require?
Answer #1

Given that t180 = 12.5 \mus = 12.5*10-6 s

The frequency of the magnetic field (\omega) = \gamma.B, where \gamma is the gyromagnetic ratio and B is the strength of the magnetic field.

The frequency of magnetic field (\omega) = 1/2t180

= 1/(2*12.5*10-6 s)

= 40 kHz

Note: 1 Hz = 1 s-1

The time that would the corresponding 90o pulse require = 12.5/2 = 6.25 \mus